a 0.252g sample pf inknown organic base is dissolved in water n titrated with a 0.14M HCl sol.   aftr d addition f 20 ml acid,a  pH of 10.7 is recorded. the equivalence pt. is reached at 40 ml HCl. acid n base combine in 1:1.find kb of base??

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Ka = [H+] * [A-] / [HA]

Rearrange to give [H+(aq)]:

[H+] = Ka * ([HA] / [A-])

Take log10 of both sides:

log[H+] = log(Ka) + log([HA] / [A-])

Multiply both sides by -1:

-log[H+] = -log(Ka) - log([HA] / [A-])

Recognizing that -log[H+] is pH, and -log(Ka) is pKa

pH = pKa - log([HA] / [A-])

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi